PLEASE ANSWER ALL

1) The pressure of the atmosphere decreases with increasing altitude in the

Choose all that apply.

mesosphere

The pressure never decreases with increasing altitude.

thermosphere

stratosphere

troposphere

2)

If you were travelling outward from the earth in a space shuttle at an altitude of 79.6 km you would be in the ______.

3a) The gas, nitrogen, is found in dry air at sea level at a concentration of 78.1 percent by volume. What is the concentration of N2 expressed in ppm? ______ppm

3b) The gas, krypton, is found in dry air at sea level at a concentration of 1.00 ppm. At a pressure of 753 mm Hg and a temperature of 27.8 °C, dry air contains about 2.46×1013 million molecules per cm3. The number of grams of krypton found in 1 cm3 of dry air at this temperature and pressure is ______grams.

3c) The gas, carbon monoxide, is found in dry air at sea level at a concentration of 1.00×10-5 percent by volume. What is the concentration of COexpressed in ppb? _____ppb
Answer #1

1. As the altitude increases the pressure decreases so if we are starting from the lowest layer i.e. troposphere the pressure will go on decreasing

2. Mesosphere as mesosphere extends from about 50 to 85 km above the earth's surface

3. % Concentration = ppm / 10,000

as one part per million means 1 in a million, so: 1 ppm = 1/106

and 1% = 1 / 100 = 0.01

so, 1% of a million = 0.01 x 1,000,000 = 10,000 ppm

(a) for 78.1 % N2

ppm = 78.1 X 10,000

Concentration of N2 = 7.81 X 105 ppm

(b) 2.46×1013 million molecules per cm3 = 2.46×1019  molecules per cm3

we are given that concentration of krypton is 1 ppm

it means in 1 million molecule one will be of krypton

so, out of 2.46×1019 molecule 2.46×1013 will be of Krypton

now, 6.022 X 1023 molecules = 1 mole

so 2.46 X  1013 will contain = [1/(6.023X1023)] X (2.46X 1013) = 0.4085 X 10-10 moles

also, number of moles = mass / molar mass

mass = (0.4085 X 10-10 ) moles X 83.798 g/mol = 34.23 X10-10 g

(c) 1 ppb = 1x10-9 = 1000 ppm and concentration of CO = 10-5 %

% Concentration = ppm / 10,000

ppm = 10-5 X 104 = 0.1

0.1 ppm = 0.1 X 10-3 = 10-4 ppb

​​​​​​​hope this helps! don't fprget to upvote:)